### v1

# Discrete Solution Operator Learning for Geometry-Dependent PDEs

Abstract
Neural operator learning accelerates PDE solution by approximating operators as mappings between continuous function spaces. Yet in many engineering settings, varying geometry induces discrete structural changes, including topological changes, abrupt changes in boundary conditions or boundary types, and changes in the effective computational domain, which break the smooth-variation premise. Here we introduce Discrete Solution Operator Learning (DiSOL), a complementary paradigm that learns discrete solution procedures rather than continuous function-space operators. DiSOL factorizes the solver into learnable stages that mirror classical discretizations: local contribution encoding, multiscale assembly, and implicit solution reconstruction on an embedded grid, thereby preserving procedure-level consistency while adapting to geometry-dependent discrete structures. Across geometry-dependent Poisson, advection–diffusion, linear elasticity, as well as spatiotemporal heat-conduction problems, DiSOL produces stable and accurate predictions under both in-distribution and strongly out-of-distribution geometries, including discontinuous boundaries and topological changes. These results highlight the need for procedural operator representations in geometry-dominated regimes and position discrete solution operator learning as a distinct, complementary direction in scientific machine learning.
Keywords Neural Operator $\cdot$ Discrete numerical methods $\cdot$ Deep Learning $\cdot$ Partial Differential Equations
1 Introduction
The numerical solution of partial differential equations (PDEs) lies at the core of scientific computing and engineering analysis [2, 26, 1, 18]. For decades, mature discrete numerical algorithms, such as finite element [26], finite volume [14], and finite difference [29] methods, have provided reliable and accurate solutions across a wide range of applications. A defining characteristic of these methods is that they operate on discrete computational representations: local contributions are evaluated on elements, assembled into global representations, and subsequently solved through well-defined procedures. While geometry variation alters the instantiated discrete system, the underlying solution procedures by which local contributions are evaluated and assembled remain unchanged (Fig. 1a). That is, the same local evaluation and assembly rules are reused wherever they are activated, even though the activated locations change with geometry. From a functional perspective, these procedures define mappings from problem specifications to solutions, but are implemented through explicit local-to-global procedures.
Recently, data-driven surrogate and operator learning methods have emerged as a promising alternative for accelerating PDE solutions across varying inputs and physical parameters [5, 24, 3, 8]. Two complementary directions have been widely explored. One treats PDE learning as field-to-field regression on discrete grids (or strong discretization-based surrogates) using convolutional architectures, often in encoder-decoder or UNet-based forms [21, 20, 35], achieving strong empirical performance when inputs and outputs are naturally represented on fixed-resolution discretizations. The other focuses on continuous neural operators—such as DeepONet [22], the Fourier Neural Operator (FNO) [17], and their variants [9, 12, 13, 15, 34, 16, 10, 23, 31, 30, 7]—that approximate smooth mappings between input and output functions in continuous function spaces (Fig. 1d). Many of these methods have demonstrated impressive performance on benchmark problems with regular domains and smoothly varying geometries [32, 11]. These advances also highlight the role of inductive bias, i.e., structural assumptions encoded in model architectures or learning objectives that shape what solution mechanisms can be learned efficiently from finite data.
However, in many practical engineering problems, geometry variation does not occur smoothly. Instead, it often manifests through sharp corners, internal holes, topological changes, or discontinuous boundary specifications [33, 28]. When geometry-induced non-smooth variations dominate the problem complexity, the mapping from inputs to solutions may not, in general, be reliably captured through approximation within a smooth function-space operator representation.
This observation highlights a distinction that is often overlooked: although classical discrete numerical solvers can be regarded as operators, learning a continuous operator is not equivalent to learning a discrete numerical solution procedure. Classical solvers do not approximate a single global mapping; instead, they execute a sequence of fixed procedures—local contribution evaluation, global assembly, and solution—which is invariant under geometry variation, even though the assembled global information changes (Figs. 1a,b) [19, 27, 6, 4]. Approximating a smooth functional mapping may therefore be misaligned with the procedural nature of geometry-dependent solution processes.
Here, we introduce Discrete Solution Operator Learning (DiSOL), a paradigm that explicitly targets operator learning towards discrete solution procedures. DiSOL shifts the learning target from approximating smooth functional mappings to learning how local contributions, boundary constraints, and multi-scale information are instantiated and assembled on discretized computational representations. Conceptually, DiSOL mirrors the canonical stages of numerical solvers, namely local evaluation, global assembly, and solution, while remaining fully data-driven and end-to-end differentiable (Fig. 1c). DiSOL should therefore be viewed not as a variant of continuous neural operators, but as a complementary paradigm with a different modeling objective and inductive bias: it preserves procedure-level invariance under geometry variation while allowing the instantiated discrete structure to adapt to geometry-dependent configurations.
We validate DiSOL on four classes of geometry-dependent PDEs with increasing complexity: a Poisson equation, an advection–diffusion equation spanning diffusion- and transport-dominated cases, a vector-valued linear elasticity problem, and a spatiotemporal thermal conduction problem. Across these cases, DiSOL delivers stable and accurate predictions under both in-distribution (ID) and out-of-distribution (OOD) computational domains, whereas continuous neural operators exhibit structural degradation when geometry-induced discrete variation dominates the problem setting. Moreover, our results indicate that these failures arise primarily from a mismatch between the assumptions of continuous operators and geometry-induced discrete solution structures, rather than from optimization difficulty or insufficient model capacity.
Taken together, our findings advocate a shift in how operator learning is conceptualized for geometry-dependent PDEs and other settings where discrete structural variation dominates the solution mechanism: learning discrete solution procedures provides a more faithful pathway toward reliable data-driven solvers that generalize across complex geometries.
Fig. 1: Discrete Solution Operator Learning (DiSOL) in contrast to continuous neural operator paradigms. a, Classical discrete numerical methods for partial differential equations (PDEs), exemplified by the finite element method. A PDE operator can be realized through a sequence of fixed discrete procedures, including local operator evaluation (e.g., element-level stiffness and force computation), global assembly, and solution via assembled global systems. Under geometry variation, the assembled results change, whereas the underlying assembly procedures remain unchanged. b, An abstracted view of discrete problem solving, highlighting the canonical workflow shared by many numerical methods: local contributions evaluation, global information assembly, and solution through globally aggregated representations. c, Discrete Solution Operator Learning (DiSOL). Instead of explicitly constructing numerical matrices, DiSOL represents the discrete solution process through learnable neural operators that conceptually correspond to the local operator, the multiscale assembly operator, and the implicit problem-solving operator. These components collectively form a discrete operator that preserves procedure-level invariance under geometry variation, in the sense that the underlying solution procedure remains unchanged, while adapting its outputs to geometry-dependent discrete structures. d, Representative continuous neural operator paradigms, illustrated by DeepONet [22] and FNO [17]. These approaches model PDE operators as global mappings in continuous function spaces, using global basis combinations or spectral transformations. Such representations are primarily designed for smooth function-space mappings, and may become misaligned when geometry-induced domain changes and discontinuities dominate the problem setting. In contrast, DiSOL does not aim to approximate continuous operators directly, but learns discrete solution procedures, leading to a fundamentally different inductive bias in geometry-dependent problem settings.
2 Results
DiSOL targets the learning of a discrete solution operator defined on a complete discrete problem specification:
$\mathcal{G}_{h}:(\Omega_{h},\Gamma_{h},f_{h},\eta_{h})\mapsto u_{h},$ (1)
where $\Omega_{h}$ denotes the discrete computational domain (geometry mask and adjacency implied by the grid), $\Gamma_{h}$ encodes boundary-type and boundary value specifications on that domain, $f_{h}$ denotes source fields, $\eta_{h}$ collects problem-dependent physical parameters (e.g., diffusion coefficient, advection velocity, elastic moduli) that are fixed or prescribed for a given task, and $u_{h}$ denotes the normalized solution pattern field on the grid. In the main text we focus on learning the normalized solution pattern $u_{h}:=U_{h}/u_{\lim}$ (see Methods, Section 4.1, and Supplementary Information D); all models predict $\hat{u}_{h}$, and all reported metrics are computed on $u_{h}$ restricted to $\Omega_{h}$. All experiments below follow Eq. 1.
Here, the subscript $h$ denotes a fixed discrete representation (grid resolution in this work) as in standard numerical discretization. All primary experiments train and evaluate at a fixed resolution, consistent with practical surrogate usage. We additionally report a preliminary zero-shot cross-resolution test in the Supplementary Information E.5. In all experiments, zero Dirichlet conditions are imposed on the selected boundary segments provided as input, while homogeneous Neumann conditions are implicitly assumed on all other boundaries, consistent with the finite-element data generation. Details of the data generation procedure and parameter settings are provided in Supplementary Information B, and detailed architecture and training settings of all models are provided in Supplementary Information C.
2.1 Geometry-dependent Poisson problem
Fig. 2: Discrete solution operator learning (DiSOL) for a geometry-dependent 2D Poisson problem. a, Problem formulation and data representation. The input to the model consists of three discrete fields defined on a fixed Cartesian grid: a geometry mask, a boundary-condition selection map, and a source term. The model outputs the corresponding discrete solution pattern. b, Training and validation loss histories for DiSOL, DeepONet and the FNO under comparable model capacity and training settings. DiSOL converges faster and achieves substantially lower validation loss (All models are trained under the same optimizer/schedule and have comparable parameter counts, $\approx 0.13$M each; details in Supplementary Information C.4 and 5.). c, ID test results for two representative cases. DiSOL accurately predicts the ground-truth solution patterns, while DeepONet shows pronounced distortions in the normalized pattern values (all fields are normalized patterns), despite the overall smoothness of the solutions. d, OOD generalization tests involving unseen geometries with sharp corners and internal holes, discontinuous and localized boundary conditions, and high-frequency source terms. DiSOL remains stable and consistent with the reference solutions, whereas DeepONet and FNO show substantial degradation in solution-pattern fidelity. See Supplementary Information A.1, B and E.1 for full setup and extended statistics.
We first consider a geometry-dependent elliptic problem as a minimal setting to isolate the effect of domain variation. In this case, the governing equation is fixed, while the computational domain and boundary configuration vary across samples. This setup allows us to examine whether an operator can generalize across discrete changes in domain topology and boundary activation, rather than merely interpolating solution patterns on a fixed geometry. The underlying equation is a scalar Poisson-type problem; detailed formulations are provided in the Supplementary Information A.1.
Unlike standard operator-learning benchmarks defined on regular domains, geometry, boundary specifications, and source distributions are explicitly provided as discrete input channels (Fig. 2a), and the solution pattern is predicted on the same computational grid.
Although both inputs and outputs are represented on a regular grid, the effective computational domain and boundary conditions may vary substantially, including the presence of internal holes, non-smooth boundaries, and highly localized boundary segments. Under this formulation, the dominant source of problem complexity no longer arises from source terms or material parameters, but instead stems from geometry-induced changes in the discrete computational structure.
The training dataset consists of topologically simple, simply connected geometries with smooth outer boundaries. Boundary conditions are applied continuously along the external boundary, and source terms are generated as linear combinations of low-frequency modes. As a result, the corresponding reference solutions are globally smooth, making this a relatively mild benchmark among geometry-dependent PDE problems.
Despite its relative simplicity, this setting already exposes a fundamental distinction between continuous and discrete operator learning paradigms. Because geometric variations directly alter the set of active computational regions and boundary segments on the grid, successful generalization requires adapting to discrete changes in local contributions rather than smooth interpolation of solution patterns.
Using identical training strategies and comparable model capacities, we train DiSOL, DeepONet, and FNO, and compare their convergence behavior and generalization performance. As shown in Fig. 2b, DiSOL converges rapidly during training and achieves substantially lower errors on both the training and validation sets. In contrast, the training error of FNO saturates at a relatively high level and does not decrease further with continued optimization. Under this setting, DeepONet attains lower validation errors than FNO, indicating relatively stronger fitting capacity. For conservative ID comparisons, we therefore adopt DeepONet as the primary continuous-operator baseline, while still reporting FNO results for completeness.
Figure 2c presents two representative ID test cases (more results are presented in Supplementary Information E.1). Although these samples match the training distribution in geometry and boundary configuration and yield globally smooth reference solutions, DeepONet already exhibits systematic deviations, most notably pattern mismatches and boundary-localized response errors near irregular geometric features. These deviations are structured rather than stochastic, indicating difficulty in capturing geometry-induced local contribution patterns within a continuous-operator parameterization. In contrast, DiSOL accurately reproduces both the global solution morphology and localized boundary responses.
We further assess generalization under strongly OOD conditions (Figs. 2d and S6), where multiple unseen factors are introduced simultaneously. These include geometries with sharp corners and internal holes, which induce topological changes and newly formed internal boundaries, spatially discontinuous and highly localized boundary conditions, and source terms containing substantial high-frequency components. Despite these challenges, DiSOL remains stable and produces solution patterns that closely match the reference solutions in both overall structure and key local features.
In contrast, DeepONet exhibits pronounced structural failure under OOD geometric conditions. Its predictions not only suffer from localized errors, but also deviate from the reference solutions at the level of global solution patterns. This behavior indicates that, for geometry-dependent discrete problems, continuous operator paradigms struggle to robustly capture the discrete structural variations induced by significant geometric changes.
Overall, even for the classical two-dimensional Poisson equation, where solution patterns are globally smooth and the underlying physical mechanisms are relatively simple, continuous neural operator methods already exhibit intrinsic limitations when the problem is formulated in a geometry-dependent and discrete manner. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL aligns more naturally with the principles of numerical solution procedures, thereby achieving robust and accurate predictions under both ID and OOD geometric conditions.
2.2 Geometry-dependent advection–diffusion problem
Fig. 3: Discrete solution operator learning (DiSOL) results for the advection-diffusion equation. a,b, Representative ID test cases under two transport cases: a diffusion-dominated case ($\text{Pe}\approx 0.45$) and an advection-dominated case ($\text{Pe}\approx 4.5$). Ground-truth solutions are compared with predictions from the proposed discrete solution operator (DiSOL) and the FNO. As transport effects become stronger and solution structures align more closely with flow direction and complex geometric boundaries, DiSOL consistently preserves the global transport patterns and boundary-induced features, whereas FNO exhibits increasingly pronounced structural deviations, particularly near irregular geometric features. c,d, Training and validation loss histories for DiSOL, FNO, and DeepONet under the same two Péclet number settings ($\text{Pe}\approx 0.45$ and $\text{Pe}\approx 4.5$). In both cases, DiSOL converges rapidly and achieves substantially lower training and validation errors, while the continuous neural operator baselines saturate at significantly higher error levels. e, OOD generalization results for the advection-diffusion equation at $\text{Pe}\approx 4.5$, involving unseen geometries with sharp corners and topological changes, together with localized and discontinuous boundary conditions. DiSOL preserves the overall transport structure and solution morphology. f, Statistical comparison of relative L1 errors across different models under ID and OOD settings for both $\text{Pe}\approx 0.45$ and $\text{Pe}\approx 4.5$. Box plots (logarithmic scale) show that DiSOL consistently outperforms continuous neural operator baselines in both cases, and that performance degradation is primarily driven by geometric distribution shifts rather than by increased transport dominance. See Supplementary Information A.2, B and E.2 for full setup and extended statistics.
We next consider a steady advection–diffusion problem to examine geometry generalization under transport-dominated dynamics. Although the governing equation and physical parameters are fixed within each case, variations in domain geometry induce sharp changes in effective transport pathways and boundary-layer structures. This case therefore probes whether an operator can adapt to geometry-induced changes in solution structure when global transport effects play a dominant role. The precise formulation and parameter settings are provided in Supplementary Information A.2.
We adopt the same discrete input–output representation as in the Poisson case (Fig. 2a), consisting of three input channels: geometry mask, boundary-condition locations, and source field; and a single output channel representing the scalar solution pattern. For each transport case, the governing transport parameters are fixed and not treated as additional inputs. The Péclet number ($\mathrm{Pe}$) characterizes the relative strength of advection to diffusion, and separate operators are trained for two representative cases ($\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$). Under this design, increased difficulty arises from the interaction between transport-driven solution structures and geometry-induced discreteness, rather than from changes in the input encoding. It is worth emphasizing that, although the input representation is consistent across different PDEs, the advection–diffusion operator is inherently more complex: its dominant solution patterns are jointly shaped by transport direction, boundary conditions, and geometric constraints, rather than being determined solely by the source distribution.
Figures 3a and b present representative ID validation results for the diffusion-dominated ($\mathrm{Pe}\approx 0.45$) and advection-dominated ($\mathrm{Pe}\approx 4.5$) cases (more results are presented in Supplementary Information E.2). At lower Péclet numbers, the solution patterns remain relatively smooth but already exhibit asymmetric structures aligned with the transport direction. In this case, DiSOL accurately reconstructs the global solution distribution and captures the dominant gradient orientations together with their interaction with geometric boundaries. In contrast, FNO struggles to maintain structural consistency on irregular geometries, producing predictions with noticeable morphological distortions across multiple regions.
As the Péclet number increases, advection becomes the dominant mechanism governing solution behavior. The resulting solution structures exhibit stronger directional transport features and heightened sensitivity to geometric boundaries and local structural variations. Under this advection-dominated setting (Fig. 3b), DiSOL continues to capture the primary transport pathways and their coupling with complex geometries, yielding solution patterns that remain consistent with the reference numerical solutions in both global morphology and key local features. By contrast, FNO displays pronounced structural distortions, with predicted patterns failing to reflect the true transport behavior, particularly near geometric boundaries and internal obstacles.
To further examine optimization behavior under varying transport conditions, Figs. 3c and d show the training and validation loss histories for the two Péclet cases. In this transport-dominated setting, FNO serves as the primary continuous-operator baseline, while DeepONet results are also reported for completeness. For both $\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$, DiSOL exhibits stable convergence and achieves substantially lower training and validation errors. In contrast, continuous neural operator baselines show early error saturation in both cases, and their relative convergence behavior does not fundamentally improve with increasing transport dominance. This observation suggests that transport strength itself is not the primary source of difficulty in operator learning.
We further evaluate generalization under strongly OOD geometric conditions at high Péclet numbers (Figs. 3e, f and S10), where geometries with sharp corners and topological changes are combined with highly localized and spatially discontinuous boundary conditions. Although these settings substantially amplify geometry-induced discrete structural variations, DiSOL maintains stable predictive performance, with error distributions that are statistically superior to those of continuous operator baselines. The corresponding boxplot analysis of relative $L_{1}$ errors indicates that performance degradation is primarily driven by geometric distribution shifts, rather than by the increase in the Péclet number itself.
Taken together, these results demonstrate that, for the advection–diffusion equation governed jointly by geometry and boundary conditions, the limitations of continuous neural operator methods do not stem from increased transport strength alone. Instead, they arise from the amplification of geometry-induced discrete structural variations under transport-dominated conditions. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL more naturally captures the coupling between geometry and transport effects, thereby achieving enhanced stability and generalization performance in both ID and OOD settings.
2.3 Geometry-dependent linear elasticity problem
Fig. 4: Discrete solution operator learning (DiSOL) results for geometry-dependent 2D linear elasticity. a, Problem formulation and data representation for the linear elastic solid mechanics problem. The input to the model consists of five discrete channels defined on a fixed Cartesian grid, encoding the geometry mask, displacement boundary conditions in the horizontal and vertical directions, and externally applied force components. The model outputs the corresponding two-component displacement field $(\hat{u}_{x},\hat{u}_{y})$ defined on the same grid. b, Representative ID test results. Ground-truth displacement fields are compared with predictions from the proposed discrete solution operator (DiSOL) for two representative geometries and loading configurations. DiSOL accurately reconstructs both horizontal and vertical displacement components, capturing the global deformation patterns and boundary-induced responses without introducing spurious oscillations or rigid-body artifacts. c, OOD generalization results involving unseen geometries with increased complexity, including internal holes, topological changes, and altered boundary condition configurations. Despite the substantial changes in geometric structure and boundary constraints, DiSOL preserves the overall deformation modes and displacement distributions, maintaining consistency with the reference solutions for both displacement components. See Supplementary Information A.3, B and E.3 for full setup and extended statistics.
Building upon the scalar-field problems discussed above, we further consider a two-dimensional linear elasticity problem to evaluate geometry generalization for vector-valued and coupled solution patterns. Under the same discrete operator formulation $\mathcal{G}_{h}$, the learning task targets a vector-valued displacement field $(u_{x},u_{y})$. The discrete problem specification is encoded through five input channels (Fig. 4a), which jointly represent the geometry mask $\Omega_{h}$ and the boundary and loading conditions required to determine the elasticity solution. Linear elasticity is a fundamental model in computational mechanics, whose solution behavior is jointly governed by geometric shape, displacement constraints, and external loads. This makes it a stringent test case for assessing whether operator learning methods can robustly handle multi-component, coupled solution patterns under geometry-dependent discrete settings. The precise formulation and parameter settings are provided in Supplementary Information A.3.
As in the previous cases, the computational domain geometry is specified by a discrete mask on a fixed Cartesian grid. The five input channels encode the geometry mask, prescribed displacement constraints in the horizontal and vertical directions, and the corresponding external force distributions in each direction (Fig. 4a). The model output is a two-component displacement field defined on the same grid, representing the horizontal ($u_{x}$) and vertical ($u_{y}$) displacement components. Compared to the scalar-field problems considered earlier, this setting introduces additional challenges at the operator level, including multi-channel inputs, vector-valued outputs, and strong coupling between boundary conditions and body forces. These factors collectively increase the structural complexity of the underlying solution operator.
Figure 4b presents representative ID test results (more results are presented in Supplementary Information E.3). DiSOL accurately reconstructs the spatial distributions of both displacement components, with predicted patterns closely matching the reference results in terms of global deformation patterns, boundary-induced displacement responses, and local gradient variations. Notably, no common non-physical artifacts, such as spurious oscillations, rigid-body drift, or inconsistencies between displacement components, are observed. This indicates that, within the training distribution, DiSOL successfully captures the discrete solution structure jointly determined by geometry, boundary conditions, and external loading.
We further evaluate generalization under OOD geometric conditions (Fig. 4c) by introducing complex geometries containing internal holes, which induce topological changes and substantially modify the discrete computational structure. Although such configurations are absent during training, DiSOL preserves the overall deformation patterns of the displacement fields. The predicted patterns remain consistent with the numerical references in their spatial organization, and no pronounced structural instability or global distortion is observed.
Taken together, these results demonstrate that the proposed discrete solution operator learning framework extends naturally beyond scalar elliptic and transport-type equations to geometry-dependent, vector-valued mechanics problems. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL captures discrete structural variations induced jointly by geometry, boundary conditions, and external loads. As a result, the framework delivers stable and reliable predictions under both ID and OOD scenarios.
2.4 Geometry-dependent thermal conduction problem
Fig. 5: Discrete solution operator learning (DiSOL) for the heat conduction problem. a, Problem formulation and data representation. The model input consists of discrete channels defined on a fixed Cartesian grid, encoding the geometry mask, boundary condition locations, source term distribution, and initial condition (IC). The time coordinate is provided as an additional input channel. The model predicts the temperature field at queried times $t$; during training, $t$ is sampled from $0,20]\,\mathrm{s}$. b, Representative ID test results at $t=2\,\mathrm{s}$, $10\,\mathrm{s}$, and $20\,\mathrm{s}$, showing close agreement between ground-truth solutions and DiSOL predictions. c, Zero-shot OOD evaluation combining geometric extrapolation and temporal forecasting. The model is tested on an unseen geometry with internal voids and altered boundary configurations. Predictions at $t=30\,\mathrm{s}$, $40\,\mathrm{s}$, and $50\,\mathrm{s}$ extend beyond the training temporal interval, corresponding to genuine future forecasting scenarios. Color bars indicate normalized temperature magnitude. See Supplementary Information A.4, B and E.4 for full setup and extended statistics.
The preceding examples have demonstrated the effectiveness of DiSOL for geometry-dependent operator learning problems in static settings. We next examine whether the proposed discrete operator learning framework extends naturally to dynamic problems involving spatiotemporal evolution. To this end, we consider a two-dimensional heat equation as a representative time-dependent PDE; the precise formulation is provided in Supplementary Information A.4.
As in the previous cases, discretized geometry, boundary conditions, source terms, and the initial condition are supplied as discrete input channels to the model (Fig. 5a). To account for temporal evolution, the time coordinate is incorporated as an additional input channel and provided in parallel with the spatial inputs. Under this formulation, the learning task corresponds to approximating a discrete spatiotemporal solution operator defined on geometry-dependent computational domains. The model is trained on trajectories spanning the temporal interval $t\in[0,20]\,\mathrm{s}$.
Figure 5b presents representative validation results at $t=2\,\mathrm{s}$, $10\,\mathrm{s}$, and $20\,\mathrm{s}$ (more results are presented in Supplementary Information E.4). DiSOL accurately reproduces the evolving temperature field across all evaluated time instances, capturing both the global diffusion patterns and geometry-induced local variations. The close agreement between predicted and reference solutions indicates stable learning of the underlying spatiotemporal dynamics within the training temporal window.
We further assess generalization under a zero-shot OOD setting (Figs. 5c and S15), where multiple unseen factors are introduced simultaneously. These include increased geometric complexity with internal voids, altered and discontinuous boundary configurations, and higher-frequency source and initial condition fields. In addition to spatial extrapolation, this test also probes temporal extrapolation: predictions at $t=30\,\mathrm{s}$, $40\,\mathrm{s}$, and $50\,\mathrm{s}$ extend beyond the training temporal interval and therefore constitute genuine future forecasting rather than interpolation. Despite this combined OOD setting in geometry, boundary conditions, source and initial fields, and time, DiSOL remains stable and produces solution patterns that closely match the numerical reference solutions across all evaluated time instances.
These results indicate that the discrete operator formulation underlying DiSOL extends coherently to dynamic, geometry-dependent problems. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, the framework preserves robustness under simultaneous spatial and temporal distribution shifts.
3 Discussion
In this work, we introduced a discrete solution operator learning (DiSOL) framework for geometry-dependent partial differential equations. Rather than approximating a global mapping between continuous function spaces, DiSOL operates directly on discretized computational domains and learns to assemble local solution contributions conditioned on geometry and boundary information. Through a series of benchmark problems—including scalar elliptic equations, transport-dominated advection–diffusion, vector-valued linear elasticity, and time-dependent heat conduction—we demonstrated that this discrete operator formulation enables robust generalization across geometry-dependent problem settings.
A central insight emerging from our results is that the primary challenge in geometry-dependent PDE learning does not stem from model capacity or optimization, but from the nature of geometric variation itself. Importantly, this behavior should not be interpreted as a fundamental limitation of continuous neural operator learning. Specifically, in many practical settings, changes in geometry induce discrete and non-smooth structural variations—such as alterations in domain topology, boundary classification, or active computational regions—that fundamentally modify how local contributions are assembled. When such geometry-induced effects dominate problem complexity, the mapping from problem specification to solution patterns is no longer well described as interpolation within a smooth function-space operator. Continuous operator methods remain highly effective when variations in problem specification, including geometry, can be treated as smooth perturbations of an underlying function-space mapping.
However, geometry-dependent problems often violate this smooth-variation premise. While the governing physical laws and numerical solution procedures remain invariant, geometric changes discretely alter how these procedures are instantiated on a computational domain. In this sense, what changes is not the underlying algorithmic logic, but the discrete structure through which local contributions are activated, combined, and propagated. Within this perspective, the advantage of DiSOL does not arise from a particular network architecture, but from its alignment with the discrete nature of geometry-dependent numerical problems. By operating directly on discretized representations, the learned operator mirrors the structure of classical numerical solvers, explicitly accommodating geometry-induced structural variation without forcing such effects into a smooth function-space interpolation framework.
This viewpoint provides a unified explanation for the consistent performance of DiSOL observed across all studied cases, ranging from scalar elliptic problems to transport-dominated dynamics, vector-valued elasticity, and spatiotemporal evolution. In each setting, performance degradation of continuous operator baselines is closely associated with geometry-induced discrete structural changes, rather than with increased physical complexity or stronger coupling alone. From a broader perspective, discrete and continuous operator learning paradigms should be viewed as complementary rather than competing: continuous operators offer elegant representations for problems dominated by smooth parametric variation, whereas discrete operator formulations are more naturally suited to cases in which geometry plays a central and structurally discrete role. The present study is limited to fixed-resolution discrete representations, and extending the framework to variable-resolution discretizations, incorporating stricter physical constraints, or developing hybrid formulations that combine discrete and continuous operator principles represent promising directions for future work.
4 Methods
4.1 Notation for normalized solutions
In this work, we let $U_{h}$ denote the ground-truth solution field with physical amplitude on the embedded grid (scalar for Poisson/advection-diffusion/heat conduction problems and vector-valued for elasticity). We define a per-sample amplitude $u_{\lim}$ (Supplementary Information. D, e.g., $u_{\lim}=\max_{x}|U_{h}(x)|$) and the dimensionless solution pattern $u_{h}:=U_{h}/u_{\lim}$, so that $\max_{x}|u_{h}(x)|=1$. All models (DiSOL and all baselines) are trained to predict the normalized pattern $\hat{u}_{h}$. If an absolute-amplitude prediction is required, it can be reconstructed as $\hat{U}_{h}=\hat{u}_{\lim}\,\hat{u}_{h}$ using an optional amplitude regressor (Supplementary Sec. D).
4.2 Problem setup and discrete learning objective
We formulate geometry-dependent PDE solving as learning a mapping from a discretized problem specification to its (discretized) solution on a fixed Cartesian grid. Each problem instance is represented by a multi-channel tensor $\boldsymbol{X}\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$ that encodes (i) a binary geometry mask $m\in\{0,1\}^{H\times W}$ indicating the computational domain $\Omega_{h}$, (ii) boundary-condition indicators on the embedded grid, and (iii) problem-dependent driving terms (e.g., sources, loads, or initial conditions). The goal is to learn an operator $\mathcal{F}_{\theta}$ that predicts the corresponding normalized solution pattern (or vector-valued patterns for elasticity) on the same grid
$\hat{u}_{h}=\mathcal{F}_{\theta}(\boldsymbol{X})\in\mathbb{R}^{C_{\text{out}}\times H\times W}.$ (2)
Within each benchmark, the operator is learned and evaluated at a fixed resolution; the grid spacing $h$ denotes the discrete representation scale rather than a refinement parameter. To enforce geometric feasibility and avoid spurious outside-domain predictions, we apply a fixed domain-projection operator $\Pi_{m}(v)=m\odot v$ ($\odot$ denotes element-wise multiplication) and compute all training losses and evaluation metrics on the projected fields (i.e., restricted to $\Omega_{h}$). This protocol ensures that comparisons are not influenced by trivial padding effects in inactive regions. Complete PDE specifications and numerical solver settings are provided in Supplementary Information A, and the construction of geometry/condition channels and dataset generation procedures are detailed in Supplementary Information B and C.1.
4.3 Discrete solution operator learning (DiSOL)
DiSOL is designed to approximate a discrete problem-solving operator by mirroring the compositional structure of classical numerical procedures on an embedded Cartesian grid. In particular, DiSOL factorizes the mapping from a discretized problem specification to its solution into three consecutive operators: (i) a local contribution operator that learns geometry- and condition-dependent local interactions, (ii) a multiscale assembly operator that aggregates these local contributions across scales to form globally consistent representations on $\Omega_{h}$, and (iii) a lightweight problem-solving operator that reconstructs the target solution pattern from the assembled representation. While DiSOL is implemented using multiscale convolutional feature hierarchies, its architectural intent is to instantiate these three discrete-procedure operators in a differentiable form, rather than to perform generic image-to-image translation.
Local contribution operator.
DiSOL applies compact convolutional operator blocks, i.e., the local operator $\mathcal{J}_{\theta}$, that act on local neighborhoods and learn stencil-like mappings from the discretized problem specification to latent local contributions. Concretely, given a multi-channel discretized specification $\boldsymbol{X}\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the local operator outputs a dense field of local contribution features
$z_{\ell}=\mathcal{J}_{\theta}(\boldsymbol{X})\in\mathbb{R}^{C_{\ell}\times H\times W},$ (3)
where each spatial location encodes geometry- and condition-dependent response information for that cell/pixel. This design emphasizes locality at the level of computation: the global solution behavior is modeled as the composition of locally supported operations whose responses adapt to the encoded geometry and boundary specifications.
Multiscale assembly operator.
The learned local contributions are then integrated by a multiscale assembly operator that propagates and fuses information from coarse to fine resolutions. Coarser representations capture long-range coupling induced by geometry and boundary constraints, whereas finer representations restore localized structures near boundaries, interfaces, and thin features. Formally, the assembly operator aggregates $z_{\ell}\in\mathbb{R}^{C_{\ell}\times H\times W}$ across scales to produce an assembled, globally consistent representation
$z_{a}=\mathcal{A}_{\theta}(z_{\ell})\in\mathbb{R}^{C_{a}\times H\times W},$ (4)
in which long-range interactions are integrated through multiresolution pathways while preserving local detail through cross-scale fusion.
Problem-solving operator.
Given the assembled multiscale representation, DiSOL applies a lightweight problem-solving operator $\mathcal{S}_{\theta}$ to map latent features to the physically meaningful output channels on the embedded grid. Specifically, it converts $z_{a}\in\mathbb{R}^{C_{a}\times H\times W}$ into the predicted normalized solution pattern
$\hat{u}_{h}=\mathcal{S}_{\theta}(z_{a})\in\mathbb{R}^{C_{\mathrm{out}}\times H\times W},$ (5)
which can be interpreted as a learned, differentiable surrogate of the final solve/reconstruction stage in a discrete procedure.
Auxiliary geometry-consistent mechanisms.
In addition to the three core operators above, DiSOL incorporates several lightweight mechanisms to improve robustness under geometry shifts. First, geometry-aware fusion is implemented via a conditioning branch that processes geometry/boundary channels to generate feature-wise modulation parameters (FiLM) [25], which scale and shift intermediate features in the main pathway, enabling geometry-dependent adaptation of local operator responses. Second, multiscale information gating injects the geometry mask into cross-scale connections to gate feature propagation and suppress out-of-domain activations, reducing spurious information leakage through inactive regions. Finally, since both inputs and outputs are represented on an embedded Cartesian grid, we apply a fixed domain-projection operator
$\Pi_{m}(v)=v\odot m.$ (6)
By doing so, all losses/metrics are computed on the projected field.
Collectively, the full DiSOL can be written as
$\hat{u}_{h}=\mathcal{F}_{\theta}(\boldsymbol{X})=\Pi_{m}\!\left[\mathcal{S}_{\theta}\circ\mathcal{A}_{\theta}\circ\mathcal{J}_{\theta}(\boldsymbol{X})\right].$ (7)
Architectural details of the local contribution operator, multiscale assembly operator, problem-solving operator, and the auxiliary FiLM/gating mechanisms are provided in Supplementary Information C.2, with model sizes reported in Supplementary Information C.5.
Although DiSOL is implemented with convolutional modules for efficiency on embedded grids, its inductive bias is procedural: it explicitly factors local contribution computation, multiscale assembly, and implicit solving, rather than performing generic image-to-image translation. A controlled U-Net baseline is evaluated under matched capacity and identical training protocols in Supplementary Information H.
4.4 Training objective and evaluation metrics
DiSOL is trained in a supervised manner using solution patterns generated by conventional numerical solvers (Supplementary Information A.5). Let $u$ denote the reference normalized solution pattern and $\hat{u}=\mathcal{F}_{\theta}(\boldsymbol{X})$ the predicted pattern on the embedded grid. To ensure that optimization and evaluation are consistent within the geometry domains, we adopt a masked training objective and compute all errors only on $\Omega_{h}$ as indicated by the geometry mask $m$. Specifically, we use a masked L1 loss
$\mathcal{L}(\theta)=\|(\hat{u}-u)\odot m\|_{1},$ (8)
where norm is computed over spatial locations (and channels when applicable). This protocol avoids trivial contributions from inactive regions and provides a consistent basis for comparing models under geometry shifts.
[…]
When absolute-amplitude solutions are required for downstream use, they can be reconstructed via an optional amplitude recovery step (Supplementary Information D), but this post-processing is not part of the main-text comparative evaluation.
Training is performed with a standard gradient-based optimizer under a unified protocol across DiSOL and all baselines. Optimization settings (optimizer, learning-rate schedule, batch size, and training epochs) are reported in Supplementary Information C.4.
5 Acknowledgements
Support from the National Natural Science Foundation of China (Grant nos. 12032014 and T2488101) is acknowledged (X.-Q. Feng). Support from the National Natural Science Foundation of China (Grant no. 12502234) is acknowledged (J. Bai).